Carlos is 7 years older than Li wei. Three years ago, the product of their ages was 294. How old is each one now?
^^ I understand how to solve it, but i just can't set the equation up.

Answer by sudhanshu_kmr(1152)


age of Li wei = x, Carlos age = x+7
three years ago, their ages are (x-3) and (x+4) respectively..
(x-3)(x+4) = 294
=> x^2 + x -12 = 294
=> x^2 + x - 306 = 0
=> x^2 +18x -17x -306 = 0
=> x(x+18) -17(x + 18) = 0
=> (x+18)(x-17)=0
so, x = 17 , x=-18 not possible.
ages are 17 and 24 .